OpenStudy (anonymous):

-7p3(4p2 + 3p - 1) a. -28p5 - 21p4 + 7p3 b. 28p6 + 21p3 - 7p c. 21p3 + 8p2 + 3p4 - 8p d. -21p5 + 8p4 - 3p3

OpenStudy (anonymous1):

Use the distributive law for multiplication over a sum. \[-7p3(4p2 + 3p - 1) = (-7p3)\times4p2 + (-7p3)\times3p -(-7p3)\times 1\] Can you continue from here?

OpenStudy (anonymous):

Uhm , Kinda

OpenStudy (anonymous1):

Multiply each term. For example, for the first term: \[(-7p^3)\times4p^2 = -28p^5\] Can you do the other multiplications similarly? You should get A as your answer.